From nch@roe.ac.uk Wed Sep 12 10:45:51 2012 Date: Fri, 7 Sep 2012 13:47:06 +0100 From: Nigel Hambly To: Eckhard Sutorius , Nicholas Cross , Rob Blake , Ross Collins , Mike Read , Clive Davenhall , Mark Holliman Cc: Lorenzo Rimoldini , Bob Mann , Jim Emerson , Keith Noddle , Norman Gray , Tom Shanks , Stephen Warren , Simon Dye , Stelios Voutsinas , Tom Kerr , Mike Irwin , Andy Lawrence at ROE , Nigel Metcalfe Subject: WFAU VDFS Science Archives project team meeting minutes, 7/9/2012 Present: MAR, ETWS, RPB, RGM, NCH Apologies: JPE, AL, STV, NG, NJC, KTN, MSH, RSC DoNM: 10am Friday October 5th 2012 in the Vista Hut (subject to non-demolition) NCH in the chair. Actions discharged this time: ----------------------------- ACTION: MAR to add notes on the VSA known issues page relating to the problems with deep stack photometry in VVV. Discharged Actions partly discharged but continuing: ----------------------------------------- ACTION: RSC to scope out the areas/situations where we'd like to test MonetDB i.e. where do the current bottlenecks lie? Continues; RGM noted that RSC was working on a "problem statement" to kick-start discussions with the MonetDB folk. Quite important to progress this one as the idea is to use up some of the remaining staff resources from the existing WFAU grant in this area. ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; (unless NJC has had the chance) ACTION: NCH to check the GCS for pawsky-related issues, and liaise with SJW over the LAS results Continues; see Survey Release below. Actions carried forward from 24/08/12 meeting --------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: Nothing new to report this time. WFCAM, VISTA and VST updates: Nothing to report this time. Comments and issues arising from CASU minutes: The team noted the minutes of the meeting of 29th August, and all the meticulous work done in tracking down some low-level bugs in one or two codes. Magically disappearing NaNs sounds like a memory leak somewhere...? Networking: ETWS noted that we're fourth in the queue to transfer the SDSS DR9 catalogue DB from Hopkins, and that if a distribution is sent to Chicago it may be quicker to get it from there. Watch this space... WSA/VSA/OSA/GES Operations: ETWS noted: -- Written some code to reinsert the deleted FITS header keywords by comparing the headers here with headers at CASU and updating them with values from the database where available, or from the CASU headers otherwise. -- Fixed the headers of 8250 VSA and 2074 WSA files. A backup of the headers is kept in /disk41/backup. Catalogue ingest for VST/ATLAS broken files is the next thing on the list of stuff to do (next week). Hardware and Systems: RPB noted: - Ramses12 has risen from the dead now that I've managed to convince windows to install on the repaired RAID1 drive. SQL Server is back and the VVV ingest database attached and consistency checked with no problems. All we need to do now is sort out replication between the VSA and the VVV and we're back to normal. - Another casualty of the Ramses12 debacle is the backup software which has now lost its database of what's backed up where. It will need to go through all the tapes and recatalogue them, which is straight forward if time consuming. Survey Data Release: RPB noted: - VHS P88 consortium release is proceeding apace and should be finished around the middle of next week. - VMC P88 consortium release should be starting on Monday. MAR noted that VIKING QC is done for the same period, so this can be placed next in the queue. Much of the meeting this time was spent discussing what to do about the somewhat buggy VVV DR1 that, at the insistence of the co-PIs, is not to be made world-public. MAR noted that the main issue is over the deep stacked K photometry problem (noted in previous minutes) which affects ~30% of the K stacks. NCH suggested that we contact the PIs and propose that an expurgated DB is created from the current one, to at least get something out in the open that is badged "VVV DR1" to go along with their release paper, and the DR1s of the other surveys. ACTION: NCH to contact VVV co-PIs to suggest expurgating the crud from VVV DR1 and making it world-public. MAR noted that communications with the ATLAS PI have produced a tentative ATLAS release schedule of a proprietary consortium release around November, to coincide with delivery of products to ESO (but it's not clear from where those products are to come - presumably CASU if it's the OB-level products, as for the earlier VISTA deliveries?). ETWS reported going through the VSA paper with a fine tooth-comb; comments from NCH and RGM are already in. Software: Nothing of note this time. Non-survey Data Release: MAR noted that a few requests have been received for VISTA non-survey PI project DB releases, so now's the time to start formalising the procedure and checking for any unforeseen issues. Astrogrid deployment, VO & Data Analysis services: ETWS asked about the UCD1+ situation and the current status of NG's work; RGM volunteered to chase this up. ACTION: RGM to chase up NG to find out what's happening with the UCD1+ Miscellaneous: RGM noted that we need to discuss our approach to the ESO surveys workshop in October - who goes, who says what etc. is still TBD. ============================================================= Nigel Hambly Tel: +44-131-668-8234 Institute for Astronomy Fax: +44-131-668-8416 School of Physics and Astronomy University of Edinburgh Email: nch@roe.ac.uk Royal Observatory Blackford Hill Edinburgh EH9 3HJ The University of Edinburgh is a charitable body, registered in Scotland, with registration number SC005336. ============================================================= -- Scanned by iCritical.